One day in math class, shelley ask the teacher, " Mr. Nelson, how old are you?"
Mr.Nelson responds: "This year Iam 3 times as ols as my sister. How ever. 6 years ago, I was five times as old as she was". How old is the math teacher?

Answer by stanbon(48546)

One day in math class, shelley ask the teacher, " Mr. Nelson, how old are you?"
Mr.Nelson responds: "This year Iam 3 times as ols as my sister. How ever. 6 years ago, I was five times as old as she was". How old is the math teacher?
"NOW" Data:
Let sister's age be "x"; then teacher age is 3x
"6 YEARS ago Data:
Sister was x-6; teacher was 3x-6
EQUATION:
"6 years ago, I was five times as old as she was"
3x-6=5(x-6)
3x-6=5x-30
2x=24
x=12
3x=36 yrs. old (age of math teacher NOW)
Cheers,
Stan H.